OpenStudy (anonymous):

write y=2/3x-1/4 in standard form with integer coefficients

OpenStudy (anonymous):

y-2/3x=-1/4 standard form is Ax+Bx=C

OpenStudy (anonymous):

3x+4y-8=0.